OpenStudy (anonymous):

passes through (6,1) and (8,-5)

OpenStudy (saifoo.khan):

equation?

OpenStudy (anonymous):

yes

hero (hero):

I'm tired of doing these

OpenStudy (saifoo.khan):

\[\Huge y = -3x+19\]

hero (hero):

I don't know how school teachers do this year after year

OpenStudy (saifoo.khan):

ask them

OpenStudy (anonymous):

is there anyway you could possibly explain that to me so i can try the other ones on my own?

hero (hero):

That's what he should have done anyway

OpenStudy (saifoo.khan):

do you know how to find the slope?

OpenStudy (anonymous):

mx+b right?

OpenStudy (saifoo.khan):

No.

OpenStudy (anonymous):

then idk what im doing

OpenStudy (saifoo.khan):

use the formula: \[slope =\frac{y_2-y_1}{x_2-x_1}\]

OpenStudy (saifoo.khan):

remember that?

OpenStudy (anonymous):

yes i do!

OpenStudy (saifoo.khan):

plug in the values to find the slope.

OpenStudy (anonymous):

soo.. my slope would be -6/2?

OpenStudy (precal):

yes your slope is -3. By the way, school teachers love to repeat themselves. Mathematics is the study of patterns.

hero (hero):

I think you mean repetition in this case

OpenStudy (saifoo.khan):

-6/2 = -3

OpenStudy (precal):

Yes, it is part of the job.

OpenStudy (anonymous):

okay, so once i find -3, what do i do?

OpenStudy (precal):

use y=mx+b m=-3 slope is -3 use (6,1) x=6 y=1 either point given will work y=mx+b 1=(-3)(6)+b 1=-18+b 19=b equation is y=-3x+19

OpenStudy (anonymous):

thank you so much!!

OpenStudy (precal):

Anytime :)